OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (anonymous):

x^2+11x+27=0

OpenStudy (anonymous):

Thanks A bunch!!

OpenStudy (anonymous):

sorry its not 11x its 12x

OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0, -3)? and okay

OpenStudy (anonymous):

Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2)?

OpenStudy (anonymous):

x^-1=0

OpenStudy (cwrw238):

(x + 3)(x + 9) = x^2 + 12x + 27 now you need to check vertex which has x coordinate of -2b/a = -b/2a = -12/2 = -6 so we have correct equation

OpenStudy (anonymous):

Which one is that?

OpenStudy (anonymous):

the one with roots -1 and 1

OpenStudy (cwrw238):

* - b/2a not -2a/b - first one

OpenStudy (anonymous):

x^2-16x+60=0

OpenStudy (anonymous):

so what are the answers for the 2 i just asked about?

OpenStudy (cwrw238):

x^2 1 = 0 is answer to second q

OpenStudy (anonymous):

x^2-16x+60=0 is the equation for roots 6 and 10

OpenStudy (anonymous):

and the first question is x^2-1=0?

OpenStudy (anonymous):

no that is your 2nd question

OpenStudy (anonymous):

x^2-1=0 equation for roots -1 and 1

OpenStudy (anonymous):

and the answer x^2-16x+60=0 is to the question Write the equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2) right?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

by simply using this formula you ca easily ger the equation x^2-sx+p=0 where s= sum of the roots and p=product of the roots

OpenStudy (anonymous):

thanks a bunch i will have some more questions in a little bit :)

OpenStudy (anonymous):

Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2, -2/3)

OpenStudy (anonymous):

x^2-x=0

OpenStudy (anonymous):

how do i show graphs? because i am stuck on a couple questions ?

OpenStudy (anonymous):

you mean you want to draw the graphs for these equations